AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 192     Entry time: Sat Feb 12 19:48:38 2022
Author: JM, CB 
Subject: Grounding and movement tests 

> Initial grounding see final grounding from elog 187. ASIC settings for all - LEC slow comparator 0x64. Note on 11/02/22 1.8*L widths with same settings were aida1-4: 300,240,250,260

ADC data items <10k

Good events fluctuate 10k-20k

1.8*L pulser widths aida1-4: 433,400,375,375(double peak)

Attachments 1-4

> New grounding implemented. 3mm grounding cable connect fee body to adaptor board. Each adaptor board is then connected directly to the cryring ground. Attachements 10-11

ADC data items <10k

Good events fluctuate 10k-20k

1.8*L pulser widths aida1-4: 340,288,288,295 (reduced double peak)

Noise in the system also seems to have decreaed in zoomed out 1.8*L plots (not attached). Aida1 still shows large noise peak channel 32000 on 1.8*L. 1.8*W no significant change, perhaps slightly reduced.

Attachements 5-9

Screencaps for following movement tests not taken as 1.8*L and 1.8*W spectra appeared the very similar as 5-9 except for slightly different widths.  

> Detectors moved in and out (pneumatic). Leakage current dropped by ~0.2 uA, torque now reads +2.8 previously was ~0. 

no change to ADC data items or good events. During movement aida3 saw momentary increase in good events to ~40k before returning to normal range. Rest of fees good events stats unaffected by movement.

1.8*L pulser widths aida1-4: 413,313,295,306

> servo motor (only) moved, torque reduced back to ~0

no change in good events stats for any fees during servo movement. 

1.8*L pulser widths aida1-4: 425,335,316,288

> Detector moved to IN position (pnuematic and servo). 19.3mm to 81.8mm on potentiometer. 

No change to good events. During movements aida3 saw momentary increase in good events to ~41k, all other fees unaffected. 

1.8*L pulser widths aida1-4: 302,221,305,293 

Noise peak at channel ~32000 on aida 1 (1.8*L) significantly reduced. little noise on other fees 

>Detector moved to fully OUT position (pneumatic and servo). 

no change in good events, during movement aida3 saw momentary increase to ~35k

1.8*L pulser widths aida1-4: 450,350,316,290

Noise peak at chn 32000 on aida 1 has increased to previous size. little noise on other fees. 

> Bias increased to 120V. Leakage current stable however slightly higher ~0.1 uA

no change in good events. 

1.8*L pulser widths aida1-4: 447,345,189,301

> Bias increased to 140V, increase in leakage current ~0.2 uA, appears stable. Increase to 150V, fluctuate in leakage current of +/-0.5uA then huge increase to >20uA trip. 

> Fee power cycle. Bias 100V, leakage current 9.9 uA ~0.1-0.2 uA higher than previous but stable. 

no change to good events

1.8*L pulser widths aida1-4: 468, 350,311, 296

> Some decrease seen with new grounding, however main source of noise still apparent and any decrease seems insignificant compared to changes due to position movement. 

> Water pipes have been zipped tied so they no loger touch motors, no change in pulser widths or noise levels however.  

Attachment 1: 2022-02-12_14-09-31aida1-2-5-1.8*W.png  197 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 2: 2022-02-12_14-08-06aida3-4-1.8*W.png  198 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 3: 2022-02-12_14-14-13aida1-2-1.8*L.png  171 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 4: 2022-02-12_14-15-32aida3-4-1.8*L.png  174 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 5: 2022-02-12_15-53-15aida1-2-1.8*W.png  200 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 6: 2022-02-12_15-51-47aida3-4-1.8*W.png  194 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 7: 2022-02-12_16-02-08aida1-2-1.8*L.png  159 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 8: 2022-02-12_16-05-04aida3-4-1.8*L.png  152 kB  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
Attachment 9: 2022-02-12_15-54-34aida1-4-rate.png  154 kB  | Hide | Hide all
2022-02-12_15-54-34aida1-4-rate.png
Attachment 10: 20220212_152215.jpg  716 kB  | Hide | Hide all
20220212_152215.jpg
Attachment 11: 20220212_171240.jpg  618 kB  | Hide | Hide all
20220212_171240.jpg
ELOG V3.1.4-unknown